I getting stumped on this problem, please help. A merchant has coffee worth $20 a pound that she wishes to mix with 80 pounds of coffee worth $90 a pound to get a mixture that can be sold for $60 a pound. How many pounds of the $20 coffee should be used?
Answer by Earlsdon(6294)

Let x = the number of pounds of the $20-coffee. After mixing this amount with the 80 pounds of $90-coffee, the merchant will have (80+x) pounds of coffee worth $60 per pound.
Let's set up the algebraic equation to solve for x.
x($20)+80($90) = (80+x)($60) Simplify and solve for x.
20x+720 = 480+60x Subtract 20x from both sides.
720 = 480+40x Subtract 480 from both sides.
240 = 40x Finally, divide both sides by 40.
60 = x
The merchant will need to mix 60 pounds of the $20-coffee with 80 pounds of the $90-coffee to obtain (80+x = 140) pounds of $60-coffee.